# Revisiting Transformers with Insights from Image Filtering

### v1

@@ III-B Distinctions of Self-Attention
$\|\bm{p}_{i}\|^{2}=\sum_{t=1}^{d/2}\left[\sin^{2}(i/T^{2t/d})+\cos^{2}(i/T^{2t/d})\right]=\frac{d}{2}.$ (20)
Then, Eqn. 19 follows simply from the fact that
@@
where $\alpha_{c}=\exp(2c^{2}+d/2)$ is a factor coming from the norms $\|\bm{y}_{i}\|=c$ and Eqn. 20. ∎
Proposition III.2 reveals at least three distinctions between the self-attention mechanism and data-dependent image filters:
@@
    The self-attention kernel redundantly (and somewhat surprisingly) possesses an extra factor of an arguably irrelevant token-to-position cross-similarity, $K_{BF}(\bm{p}_{i},\bm{y}_{j},\bm{p}_{j},\bm{y}_{i})$, as in Eqn. 19.
@@
$K_{BSA}(\bm{p}_{i},\bm{p}_{j},\bm{y}_{i},\bm{y}_{j}):=\exp\left(\frac{\bm{p}_{i}^{\top}\bm{W}\bm{p}_{j}}{h_{p}^{2}}+\frac{\bm{y}_{i}^{\top}\bm{W}\bm{y}_{j}}{h_{y}^{2}}\right),$ (21)
where $\bm{W}=\bm{W}_{Q}^{\top}\bm{W}_{K}$. In addition, we discuss an existing attention variant from prior work that disentangles spatial and token-wise distances by introducing separate scaling parameters, $h_{p}$ and $h_{y}$ [30]. While it has already been empirically studied, our framework provides natural theoretical justification for this form of decoupled modeling. When $h_{p}^{2}\gg h_{y}^{2}$ (or $h_{p}\to\infty$ in theory), the weight values given by Eqn. 21 essentially becomes virtually insensitive to the geometric closeness of the tokens; thus, we refer to such attention as Nonlocal Self-Attention (NLSA), borrowing the terminology from the NLM filter in image processing. In practice, it would be equivalent to Transformer with no positional encoding [31]. Note that one could precompute the positional term in Eqn. 21 for all layers to eliminate all extra computational costs during inference since they are input independent.
Figure 1: Left figure: blue circles indicate the estimated tight Lipschitz constant of $\mathrm{softmax}$ as in Remark III.5 while the red curve corresponds to the best fit curve of the form $\frac{a}{\sqrt{N}}+b$. Right figure: empirical estimate of the expectation mentioned in Theorem III.2 using a random source vector $\bm{c}\in\mathbb{R}^{N}$ with $\eta_{i}\sim\mathcal{N}(0,1)$, for input dimensionality $10^{2}\leq N\leq 10^{4}$ using $10^{3}$ random samples. One can confirm that the perturbation magnitude of attention scores does not necessarily vanish with large $N$, supporting the claim of Theorem III.2.
Figure 2: Left to Right: Token-to-Token, Token-to-Position, Position-to-Token, and Position-to-Position Correlation Logit Matrices [30].
Remark III.4 (Bilateral $\sim$ Relative PE).
Note that Eqn. 21 can be seen as concatenating absolute PE with input tokens instead of adding, and is akin to relative PE where the relative distance is captured by the weighted dot product of the correcponding positional vectors. In fact, any distance proxy [^1] We use the term proxy since $d_{\theta}(i,j)$ does not have to be a proper distance metric. It is only required to decay as relative distance increases. $d_{\theta}(i,j)$ can effectively be used to implement this bilateral framework. For example, while $d_{\theta}(i,j)=\bm{p}_{i}^{\top}\bm{W}\bm{p}_{j}$ recovers the one in Eqn. 21, $d_{\theta}(i,j)=-m\left|i-j\right|$ recovers ALiBi [21] – a relative PE instance proposed for a better long sequence modeling. For completeness, we compare both variants across different backbone models and various tasks in Table I.
The kernel function defined in Eqn. 21 is similar to the kernels used by [8] when the weight matrix is set to be symmetric positive definite and [30] when positional vectors and word embeddings are equipped with independent weight matrices. The positive results reported in those works also support the idea of purely bilateral self-attention derived from our general framework.
@@
III-C Long Range Unstability Analysis of Standard Self-Attention
It has been observed empirically that token-to-position cross dot-product does not capture any particular pattern but is distributed uniformly [30], effectively acting as noise. Since both token and position vectors are $d$-dimensional independent vectors with large $d$, they are close to orthogonal in expectation [32] under mild distributional assumptions. Thus, the assumption
@@ III-C Long Range Unstability Analysis of Standard Self-Attention
$\|\mathrm{softmax}(\bm{x})-\mathrm{softmax}(\bm{y})\|\leq\lambda\|\bm{x}-\bm{y}\|,$ (22)
@@
Theorem III.2 (Impact on Attention Scores).
@@
$\mathbb{E}_{\bm{\eta}\sim\mathcal{D}(\bm{0},\sigma^{2}\bm{I})}\left[\|\mathrm{softmax}(\bm{c}+\bm{\eta})-\mathrm{softmax}(\bm{c})\|\right]\leq\sigma L\sqrt{N},$ (23)
@@
Remark III.5.
Although $\mathrm{softmax}$ is globally Lipschitz as demonstrated in Proposition III.3, local behavior can be slightly more complex also depending on input dimension $N$. As depicted in Figure 1, the optimal Lipschitz constant of $\mathrm{softmax}$, defined as
$L:=\sup_{\bm{x},\bm{y}\in S,\bm{x}\neq\bm{y}}\frac{\|\mathrm{softmax}(\bm{x})-\mathrm{softmax}(\bm{y})\|}{\|\bm{x}-\bm{y}\|},$ (24)
can behave (e.g. when a few components dominate) similarly to $1/\sqrt{N}$. Hence, while the upper bound in Theorem III.2 seems to be unbounded, the dependence of $L$ on $N$ can make it locally $O(1)$. We can, therefore, obtain the following attainable asymptotical upper bound:
$\mathbb{E}_{\bm{\eta}\sim\mathcal{D}(\bm{0},\sigma^{2}\bm{I})}\left[\|\mathrm{softmax}(\bm{c}+\bm{\eta})-\mathrm{softmax}(\bm{c})\|\right]\lesssim\sigma.$ (25)
Figure 3: LRA learning curves of baseline [1] and bilateral attention mechanisms for Image (1K) and Text (4K) datasets. The figures show that the vanilla self-attention either learns substantially slower compared to its bilateral counterpart or struggles to learn useful long context relations.
The following immediate corollary of Theorem III.2 demonstrates the effect of perturbation directly on the self-attention output vectors:
Corollary III.2 (Impact on Attention Layer Output).
@@
$\mathbb{E}_{\bm{\eta}\sim\mathcal{D}(\bm{0},\sigma^{2}\bm{I})}$ $\left[\|\mathrm{softmax}(\bm{c}+\bm{\eta})\bm{V}-\mathrm{softmax}(\bm{c})\bm{V}\|\right]$
$\quad\leq\sigma L\left\|\bm{V}\right\|_{\mathrm{op}}\sqrt{N}\asymp\sigma LN\sqrt{d},$
@@
Corollary III.2 provides the corresponding error bound on each output vector of self-attention layer since each component inside the difference norm is equivalent to one row of self-attention layer output matrix. The fact that $\|\bm{V}\|\asymp\sqrt{dN}$ (can be proved using the technique in the proof of Theorem III.2 in Appendix VII) is generally an increasing quantity as $N\to\infty$ implies it contributes a notable factor to the upper bound, potentially distracting the model from learning stable representations over a sequence of considerable length. We shall provide empirical support for this claim in Section V via long sequence modeling.
IV Rethinking Residual Connection with Boosting
Although residual connections help to increase the signal magnitude at each step, they cannot completely prevent it from vanishing as mentioned in Remark III.3. To solve this issue, we will discuss leveraging older residuals aligning with a more sophisticated boosting techniques in iterative image processing.
Definition IV.1 (Generalized Residual Connection).
@@ IV Rethinking Residual Connection with Boosting
Proposition IV.1 shows that to tackle the issue of information vanishing, one ought to bound the index sequence $\{i_{\ell}\}_{\ell\in\mathbb{N}}$ for the GRC. A reasonable GRC would be to have a constant sequence $i_{\ell}=k\in\mathbb{N}$ for all $\ell>k$. To this end, we propose the boosting framework, inspired by a boosting technique for iterative image filtering [28]. With boosting, the output of self-attention is computed as
$\bm{Y}_{\ell+1}=f_{\ell}(\bm{Y}_{\ell})+t\bm{Y}_{0}+(1-t)\bm{Y}_{\ell},$ (26)
where $t$ is a scale parameter. Rolling the Eqn. 26 for one more layer (without MLP and layer normalization for simplicity), the next transformer layer can be represented as
@@
$=f_{\ell+1}(f_{\ell}(\bm{Y}_{\ell})+\bm{Y}_{\ell})+tf_{\ell+1}(\bm{Y}_{0}-\bm{Y}_{\ell})+\bm{R}_{\ell+1}.$ (27)
The boosting procedure, therefore, can essentially be seen as strengthening the signal by overlaying the residual and current signals (higher SNR due to Proposition III.1), filtering noise in a strengthened signal. Noting that the attention computation is a pseudo-linear (linear with data-dependent entries) transformation, Eqn. 27 can recover many boosting techniques for image denoising algorithms such as twicing. Twicing, originally proposed by [34], effectively re-uses the smoothed residual after each iteration step of denoising by combining it with the new estimate to get a more fine-grained output. In Figure 5, we verify that vanilla attention with the standard residual connection tends to progressively smooth the input tokens across layers, and boosting alleviates the issue.
@@
$\bm{Y}_{\ell+1}=f_{\ell}(\bm{Y}_{\ell})+\bm{Y}_{\ell},$ (28)
to the modified, generalized residual connection (GRC) as given by Eqn. 26, where $f_{\ell}(\cdot)$ represents the layer computation (e.g., self-attention), and $t\in(0,1]$ is a parameter. The Lipschitz constant of a function quantifies its sensitivity to input changes. A function $g$ is $K$-Lipschitz if for all $\bm{x},\bm{x}^{\prime}$, $\|g(\bm{x})-g(\bm{x}^{\prime})\|\leq K\|\bm{x}-\bm{x}^{\prime}\|$. We define $K_{\ell}$ such that $\|\bm{Y}_{\ell}-\bm{Y}_{\ell}^{\prime}\|\leq K_{\ell}\|\bm{Y}_{0}-\bm{Y}_{0}^{\prime}\|$, and aim to compare $K_{n}$, the $n$-layer Lipschitzness:
Theorem IV.1 (Robustness of Boosting).
@@
$\frac{K^{\text{GRC}}_{n}}{K^{\text{RC}}_{n}}\asymp\left(1-\frac{t}{L+1}\right)^{n},$ (29)
@@
V Experimental Results
V-A Language Modeling on WikiText-103
@@
We follow the experimental setup of [25] by training a small backbone that consists of 16 layers, 8 attention heads with a dimension of 16, a feedforward layer with a size of 2048, and an embedding dimension of 128. The results in Table II indicate a general trend of improvement in language modeling performance under both clean and contaminated data settings on the WikiText-103 dataset. Starting with the baseline Transformer model, we observe a decrease in perplexity (PPL) when a boosting mechanism and Bilateral Attention is added. The contamination attack follows the method of random word swapping as described in [35].
V-B Long Sequence Modeling on the LRA Benchmark
@@
To empirically verify the propositions made in Section III-C, we experiment on long sequence modeling, adopting the setup of [36]. For each of the 5 tasks in the Long Range Arena (LRA) benchmark [37], equation calculation (ListOps), review classification (Text), document retrieval (Retrieval), image classification (Image), and image spatial dependencies (Pathfinder), we compare Bilateral Attention against standard self-attention [1] as well as ALiBi [21]. Model backbones for all three models–Transformer [1], Linformer [38], and Nyströmformer [39]–are set with 2 layers, hidden dimension of 128, 2 attention heads of dimension 32, and embedding dimension of 64. Only for the case of Linformer, we interpolate between standard PE [1] and other two models by combining input tokens with absolute PE vectors, scaled down by 0.3, before integrating ALiBi or Bilateral Attention. This introduces controlled noise in the attention scores to mitigate rapid overfitting due to the considerable number of new learnable parameters introduced by Linformer. All other training configurations are kept the same for maximum fair comparison across the variants.
@@
It is observed that models with Bilateral Attention outperform baseline models and ALiBi in 4 of 5 tasks (Table I), especially with significantly faster convergence than absolute PE during training as shown in Figure 3. This is particularly evident from the hardest task in the benchmark–document retrieval (Retrieval). The context length for the retrieval task is 4K and the required attention span (RAS) is around 1.3K, making it the most demanding among all reported tasks.
Table I: Test accuracy comparison on the Long Range Arena (LRA) benchmark for vanilla Transformer [1], Linformer [38], and Nyströmformer [39] with absolute positional encodings [1] and ALiBi [21], compared to Bilateral Attention. We use Bilateral and ALiBi to denote the two models from Remark III.4.
Dataset (seq. len.) RAS [^2] Given the purpose of LRA, required attention span (RAS) serves as a proxy for how difficult a task is for Transformer-based models. The numbers are approximate copies of Figure 2 in [37]. Trans. +ALiBi +Bilateral Lin. +ALiBi +Bilateral Nyström. +ALiBi +Bilateral
Pathfinder (1K) 600 67.83 62.31 62.27 51.22 60.82 51.22 68.22 61.07 69.00
Image (1K) 400 39.90 43.42 39.92 42.69 41.64 42.80 41.23 41.27 45.56
ListOps (2K) 800 18.55 18.35 37.70 19.15 37.15 38.26 18.70 20.82 37.00
Text (4K) 300 62.73 72.95 66.85 55.94 55.37 57.05 62.16 67.48 66.18
Retrieval (4K) 1300 61.31 74.47 76.38 77.68 78.58 78.65 60.89 70.11 73.84
Table II: Language Modeling on Wikitext-103
@@
V-C Vision Tasks on ImageNet-1k and ADE20k
In Table III and Table IV, we report the results on image classification on ImageNet-1k [40] and segmentation on ADE20k [41], comparing 4 models: DeiT [6] as a baseline, DeiT+Boost is our integration of GRC on top of residual connection, DeiT+Bilateral is our implementation of Bilateral Attention within the DeiT model replacing the standard self-attention, and DeiT+Bilateral+Boost is the model for integrating GRC on top of Bilateral Attention. The DeiT backbone uses 12 layers, 3 heads of dimension 64, patch size 16, feedforward layer of size 768 and embedding dimension of 192 similar to [25], trained for 300 epochs. For DeiT+Boost, we set the parameter $t$ to learnable with initial value 00. The experiments are carried on clean images as well as under two widely adopted adversarial attacks Fast Gradient Sign Method (FGSM) [42] and Projected Gradient Descent (PGD) [43] attacks. Additionally, Fully Attentional Network (FAN) backbone [44] is also tested in Table III together with our Boosting method to verify its ability to improve robustness of a robust backbone. Both FAN and and FAN+Boost models are trained for 130 epochs, adopting the training settings from [25].
Table III: Object Classification on ImageNet-1K without and with adversarial attack of perturbation budget 1/255.

### v2

@@ 3.2 Distinctions of Self-Attention
(23) $\|\bm{p}_{i}\|^{2}=\sum_{t=1}^{d/2}\left[\sin^{2}(i/T^{2t/d})+\cos^{2}(i/T^{2t/d})\right]=\frac{d}{2}.$
Then, Eqn. 22 follows simply from the fact that
@@
where $\alpha_{c}=\exp(2c^{2}+d/2)$ is a factor coming from the norms $\|\bm{y}_{i}\|=c$ and Eqn. 23.
Proposition 3.2 reveals at least three distinctions between the self-attention mechanism and data-dependent image filters:
@@
    The self-attention kernel redundantly (and somewhat surprisingly) possesses an extra factor of an arguably irrelevant token-to-position cross-similarity, $K_{BF}(\bm{p}_{i},\bm{y}_{j},\bm{p}_{j},\bm{y}_{i})$, as in Eqn. 22.
@@
(24) $K_{BSA}(\bm{p}_{i},\bm{p}_{j},\bm{y}_{i},\bm{y}_{j}):=\exp\left(\frac{\bm{p}_{i}^{\top}\bm{W}\bm{p}_{j}}{h_{p}^{2}}+\frac{\bm{y}_{i}^{\top}\bm{W}\bm{y}_{j}}{h_{y}^{2}}\right),$
where $\bm{W}=\bm{W}_{Q}^{\top}\bm{W}_{K}$. In addition, we discuss an existing attention variant from prior work that disentangles spatial and token-wise distances by introducing separate scaling parameters, $h_{p}$ and $h_{y}$ [ke2021rethinking]. While it has already been empirically studied, our framework provides natural theoretical justification for this form of decoupled modeling. When $h_{p}^{2}\gg h_{y}^{2}$ (or $h_{p}\to\infty$ in theory), the weight values given by Eqn. 24 essentially becomes virtually insensitive to the geometric closeness of the tokens; thus, we refer to such attention as Nonlocal Self-Attention (NLSA), borrowing the terminology from the NLM filter in image processing. In practice, it would be equivalent to Transformer with no positional encoding [haviv2022transformer]. Note that one could precompute the positional term in Eqn. 24 for all layers to eliminate all extra computational costs during inference since they are input independent.
Figure 1: Left to Right: Token-to-Token, Token-to-Position, Position-to-Token, and Position-to-Position Correlation Logit Matrices [ke2021rethinking].
Remark 3.4 (Bilateral $\sim$ Relative PE).
Note that Eqn. 24 can be seen as concatenating absolute PE with input tokens instead of adding, and is akin to relative PE where the relative distance is captured by the weighted dot product of the correcponding positional vectors. In fact, any distance proxy [^1] We use the term proxy since $d_{\theta}(i,j)$ does not have to be a proper distance metric. It is only required to decay as relative distance increases. $d_{\theta}(i,j)$ can effectively be used to implement this bilateral framework. For example, while $d_{\theta}(i,j)=\bm{p}_{i}^{\top}\bm{W}\bm{p}_{j}$ recovers the one in Eqn. 24, $d_{\theta}(i,j)=-m\left|i-j\right|$ recovers ALiBi [press2022train] – a relative PE instance proposed for a better long sequence modeling. For completeness, we compare both variants across different backbone models and various tasks in Table 1.
The kernel function defined in Eqn. 24 is similar to the kernels used by [tsai2019transformer] when the weight matrix is set to be symmetric positive definite and [ke2021rethinking] when positional vectors and word embeddings are equipped with independent weight matrices. The positive results reported in those works also support the idea of purely bilateral self-attention derived from our general framework.
@@
Figure 2: Left figure: blue circles indicate the estimated tight Lipschitz constant of $\mathrm{softmax}$ as in Remark 3.5 while the red curve corresponds to the best fit curve of the form $\frac{a}{\sqrt{N}}+b$. Right figure: empirical estimate of the expectation mentioned in Theorem 3.2 using a random source vector $\bm{c}\in\mathbb{R}^{N}$ with $\eta_{i}\sim\mathcal{N}(0,1)$, for input dimensionality $10^{2}\leq N\leq 10^{4}$ using $10^{3}$ random samples. One can confirm that the perturbation magnitude of attention scores does not necessarily vanish with large $N$, supporting the claim of Theorem 3.2.
3.3 Long Range Unstability Analysis of Standard Self-Attention
It has been observed empirically that token-to-position cross dot-product does not capture any particular pattern but is distributed uniformly [ke2021rethinking], effectively acting as noise. Since both token and position vectors are $d$-dimensional independent vectors with large $d$, they are close to orthogonal in expectation [vershynin_high-dimensional_2018] under mild distributional assumptions. Thus, the assumption
@@ 3.3 Long Range Unstability Analysis of Standard Self-Attention
(25) $\|\mathrm{softmax}(\bm{x})-\mathrm{softmax}(\bm{y})\|\leq\lambda\|\bm{x}-\bm{y}\|,$
@@
Theorem 3.2 (Impact on Attention Scores).
@@
(26) $\mathbb{E}_{\bm{\eta}\sim\mathcal{D}(\bm{0},\sigma^{2}\bm{I})}\left[\|\mathrm{softmax}(\bm{c}+\bm{\eta})-\mathrm{softmax}(\bm{c})\|\right]\leq\sigma L\sqrt{N},$
@@
Remark 3.5.
Although $\mathrm{softmax}$ is globally Lipschitz as demonstrated in Proposition 3.3, local behavior can be slightly more complex also depending on input dimension $N$. As depicted in Figure 2, the optimal Lipschitz constant of $\mathrm{softmax}$, defined as
(27) $L:=\sup_{\bm{x},\bm{y}\in S,\bm{x}\neq\bm{y}}\frac{\|\mathrm{softmax}(\bm{x})-\mathrm{softmax}(\bm{y})\|}{\|\bm{x}-\bm{y}\|},$
can behave (e.g. when a few components dominate) similarly to $1/\sqrt{N}$. Hence, while the upper bound in Theorem 3.2 seems to be unbounded, the dependence of $L$ on $N$ can make it locally $O(1)$. We can, therefore, obtain the following attainable asymptotical upper bound:
(28) $\mathbb{E}_{\bm{\eta}\sim\mathcal{D}(\bm{0},\sigma^{2}\bm{I})}\left[\|\mathrm{softmax}(\bm{c}+\bm{\eta})-\mathrm{softmax}(\bm{c})\|\right]\lesssim\sigma.$
The proof of Theorem 3.2 is deferred to Appendix. The following immediate corollary of Theorem 3.2 demonstrates the effect of perturbation directly on the self-attention output vectors:
Corollary 3.2 (Impact on Attention Layer Output).
@@
$\mathbb{E}_{\bm{\eta}\sim\mathcal{D}(\bm{0},\sigma^{2}\bm{I})}\left[\|\mathrm{softmax}(\bm{c}+\bm{\eta})\bm{V}-\mathrm{softmax}(\bm{c})\bm{V}\|\right]$ $\leq\sigma L\left\|\bm{V}\right\|_{\mathrm{op}}\sqrt{N}$
$\asymp\sigma LN\sqrt{d},$
@@
Corollary 3.2 provides the corresponding error bound on each output vector of self-attention layer since each component inside the difference norm is equivalent to one row of self-attention layer output matrix. The fact that $\|\bm{V}\|\asymp\sqrt{dN}$ (can be proved using the technique in the proof of Theorem 3.2 in Appendix) is generally an increasing quantity as $N\to\infty$ implies it contributes a notable factor to the upper bound, potentially distracting the model from learning stable representations over a sequence of considerable length. We shall provide empirical support for this claim in Section 5 via long sequence modeling.
4 Rethinking Residual Connection with Boosting
Although residual connections help to increase the signal magnitude at each step, they cannot completely prevent it from vanishing as mentioned in Remark 3.3. To solve this issue, we will discuss leveraging older residuals aligning with a more sophisticated boosting techniques in iterative image processing.
Definition 4.1 (Generalized Residual Connection).
@@ 4 Rethinking Residual Connection with Boosting
Proof. Suppose the assumptions described in Remark 3.2 hold. Then, the amount of salient information captured by $\bm{y}_{\ell+1}$ is given by
(29) $s_{\ell}:=\alpha^{\ell}\|\bm{u}_{0}\|+\alpha^{i_{\ell}}\|\bm{u}_{0}\|=\alpha^{i_{\ell}}(\alpha^{\ell-i_{\ell}}+1)\|\bm{u}_{0}\|.$
(30) $\lim_{\ell\to\infty}s_{\ell}=\lim_{\ell\to\infty}\alpha^{i_{\ell}}(\alpha^{\ell-i_{\ell}}+1)\|\bm{u}_{0}\|=0.$
Therefore, it is necessary for the sequence $\{i_{\ell}\}_{\ell\in\mathbb{N}}$ to contain at least one bounded subsequence to avoid token uniformity.
Proposition 4.1 shows that to tackle the issue of information vanishing, one ought to bound the index sequence $\{i_{\ell}\}_{\ell\in\mathbb{N}}$ for the GRC. A reasonable GRC would be to have a constant sequence $i_{\ell}=k\in\mathbb{N}$ for all $\ell>k$. To this end, we propose the boosting framework, inspired by a boosting technique for iterative image filtering [romano2015boosting]. With boosting, the output of self-attention is computed as
(31) $\bm{Y}_{\ell+1}=f_{\ell}(\bm{Y}_{\ell})+t\bm{Y}_{0}+(1-t)\bm{Y}_{\ell},$
where $t$ is a scale parameter. Rolling the Eqn. 31 for one more layer (without MLP and layer normalization for simplicity), the next transformer layer can be represented as
@@
(32) $=f_{\ell+1}(f_{\ell}(\bm{Y}_{\ell})+\bm{Y}_{\ell})+tf_{\ell+1}(\bm{Y}_{0}-\bm{Y}_{\ell})+\bm{R}_{\ell+1}.$
The boosting procedure, therefore, can essentially be seen as strengthening the signal by overlaying the residual and current signals (higher SNR due to Proposition 3.1), filtering noise in a strengthened signal. Noting that the attention computation is a pseudo-linear (linear with data-dependent entries) transformation, Eqn. 32 can recover many boosting techniques for image denoising algorithms such as twicing. Twicing, originally proposed by [tukey77], effectively re-uses the smoothed residual after each iteration step of denoising by combining it with the new estimate to get a more fine-grained output. In Figure 6, we verify that vanilla attention with the standard residual connection tends to progressively smooth the input tokens across layers, and boosting alleviates the issue.
@@
(33) $\bm{Y}_{\ell+1}=f_{\ell}(\bm{Y}_{\ell})+\bm{Y}_{\ell},$
to the modified, generalized residual connection (GRC) as given by Eqn. 31, where $f_{\ell}(\cdot)$ represents the layer computation (e.g., self-attention), and $t\in(0,1]$ is a parameter. The Lipschitz constant of a function quantifies its sensitivity to input changes. A function $g$ is $K$-Lipschitz if for all $\bm{x},\bm{x}^{\prime}$, $\|g(\bm{x})-g(\bm{x}^{\prime})\|\leq K\|\bm{x}-\bm{x}^{\prime}\|$. We define $K_{\ell}$ such that $\|\bm{Y}_{\ell}-\bm{Y}_{\ell}^{\prime}\|\leq K_{\ell}\|\bm{Y}_{0}-\bm{Y}_{0}^{\prime}\|$, and aim to compare $K_{n}$, the $n$-layer Lipschitzness:
Theorem 4.1 (Robustness of Boosting).
@@
(34) $\frac{K^{\text{GRC}}_{n}}{K^{\text{RC}}_{n}}\asymp\left(1-\frac{t}{L+1}\right)^{n},$
@@
The proof of Theorem 4.1 is provided in Appendix.
5 Experimental Results
5.1 Language Modeling on WikiText-103
@@
We follow the experimental setup of [nielsen2024elliptical] by training a small backbone that consists of 16 layers, 8 attention heads with a dimension of 16, a feedforward layer with a size of 2048, and an embedding dimension of 128. The results in Table 2 indicate a general trend of improvement in language modeling performance under both clean and contaminated data settings on the WikiText-103 dataset. Starting with the baseline Transformer model, we observe a decrease in perplexity (PPL) when a boosting mechanism and Bilateral Attention is added. The contamination attack follows the method of random word swapping as described in [merity2016pointer].
5.2 Long Sequence Modeling on the LRA Benchmark
@@
To empirically verify the propositions made in Section 3.3, we experiment on long sequence modeling, adopting the setup of [chen2023primal]. For each of the 5 tasks in the Long Range Arena (LRA) benchmark [tay2020long], equation calculation (ListOps), review classification (Text), document retrieval (Retrieval), image classification (Image), and image spatial dependencies (Pathfinder), we compare Bilateral Attention against standard self-attention [vaswani2017attention] as well as ALiBi [press2022train]. Model backbones for all three models–Transformer [vaswani2017attention], Linformer [wang2020linformer], and Nyströmformer [xiong2021nystromformer]–are set with 2 layers, hidden dimension of 128, 2 attention heads of dimension 32, and embedding dimension of 64. Only for the case of Linformer, we interpolate between standard PE [vaswani2017attention] and other two models by combining input tokens with absolute PE vectors, scaled down by 0.3, before integrating ALiBi or Bilateral Attention. This introduces controlled noise in the attention scores to mitigate rapid overfitting due to the considerable number of new learnable parameters introduced by Linformer. All other training configurations are kept the same for maximum fair comparison across the variants.
@@
It is observed that models with Bilateral Attention outperform baseline models and ALiBi in 4 of 5 tasks (Table 1), especially with significantly faster convergence than absolute PE during training as shown in Figure 3. This is particularly evident from the hardest task in the benchmark–document retrieval (Retrieval). The context length for the retrieval task is 4K and the required attention span (RAS) is around 1.3K, making it the most demanding among all reported tasks.
Figure 3: LRA [tay2020long] learning curves of baseline [vaswani2017attention] and bilateral attention mechanisms for Image (1K) and Text (4K) datasets. The figures show that the vanilla self-attention either learns substantially slower compared to its bilateral counterpart or struggles to learn useful long context relations.
Table 1: Test accuracy comparison on the Long Range Arena (LRA) benchmark, transposed so that methods are rows and tasks are columns.
Method Pathfinder (1K) Image (1K) ListOps (2K) Text (4K) Retrieval (4K)
Transformer [vaswani2017attention] 67.83 39.90 18.55 62.73 61.31
+ ALiBi 62.31 43.42 18.35 72.95 74.47
+ Bilateral 62.27 39.92 37.70 66.85 76.38
Linformer [wang2020linformer] 51.22 42.69 19.15 55.94 77.68
+ ALiBi 60.82 41.64 37.15 55.37 78.58
+ Bilateral 51.22 42.80 38.26 57.05 78.65
Nyströmformer [xiong2021nystromformer] 68.22 41.23 18.70 62.16 60.89
+ ALiBi 61.07 41.27 20.82 67.48 70.11
+ Bilateral 69.00 45.56 37.00 66.18 73.84
Table 2: Language Modeling on Wikitext-103
@@
5.3 Vision Tasks on ImageNet-1k and ADE20k
In Table 3 and Table 4, we report the results on image classification on ImageNet-1k [russakovsky2015imagenet] and segmentation on ADE20k [zhou2019semantic], comparing 4 models: DeiT [touvron2021training] as a baseline, DeiT+Boost is our integration of GRC on top of residual connection, DeiT+Bilateral is our implementation of Bilateral Attention within the DeiT model replacing the standard self-attention, and DeiT+Bilateral+Boost is the model for integrating GRC on top of Bilateral Attention. The DeiT backbone uses 12 layers, 3 heads of dimension 64, patch size 16, feedforward layer of size 768 and embedding dimension of 192 similar to [nielsen2024elliptical], trained for 300 epochs. For DeiT+Boost, we set the parameter $t$ to learnable with initial value 0. The experiments are carried on clean images as well as under two widely adopted adversarial attacks Fast Gradient Sign Method (FGSM) [madry2017towards] and Projected Gradient Descent (PGD) [goodfellow2014explaining] attacks. Additionally, Fully Attentional Network (FAN) backbone [zhou2022understanding] is also tested in Table 3 together with our lightweight Boosting method to verify its compatibility with different backbones. Both FAN and and FAN+Boost models are trained for 130 epochs, adopting the training settings from [nielsen2024elliptical].
Table 3: Object Classification on ImageNet-1K without and with adversarial attack of perturbation budget 1/255.